4) 8) $25,000 is invested at 6% compounded quarterly how much money is in the account after 10 year?

can you please hand write the steps and the answer please. i need the answer please
Answer #1

Given that,

Amount invested PV = $25000

interest rate r = 6% per year compounded quarterly

So, amount after t = 10 years can be calculated using compounding formula

Amount after 10 year FV = PV*(1 + r/n)^(n*t)

where n is compounding frequency

here n = 4 quarters in a year

So, FV = 25000*(1 + 0.06/4)^(4*10) = $45350.46

So, money in the account after 10 years = $45350.46